OpenStudy (anonymous):

what is the laplace inverse of (s^4+6s-18)/(s^5-3s^4)?

OpenStudy (anonymous):

use partial fractions and take s^4 common from denomiator it will be easy to solve

OpenStudy (anonymous):

and then use the formulae that laplace inverse of 1/s = 1 and laplace inverse of 1/s^2= t etc and laplace inverse of 1/s-3= \[e^{3t}\]

OpenStudy (anonymous):

the answer is (e^(3t) )+t^3

OpenStudy (anonymous):

would you please give me the solved question

OpenStudy (anonymous):

do you know how to make partial fractions??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then use partial fractions and it will be easy use the formulae i mentiond above

OpenStudy (anonymous):

is it form bsc?? methods

OpenStudy (anonymous):

s^4+6s-18=A/s+B/s^2+C/s^3+D/s^4+E/s-3

OpenStudy (anonymous):

yes right and then find A,B,C,D,E AND IT WILL BE 101 PERCENT SIMPLE

OpenStudy (anonymous):

ok thanks alot